the length of the rectangle is 2m less than twice the width. if the area of the rectangle is 112m^2, what is the length?
Answer by ReadingBoosters(3246)

l = 2w - 2
a = 112 = lw
112=(2w-2)(w)
112=2w^2-2w
2w^2 - 2w - 112 = 0
2(w^2 - w - 56) = 0
2(w+ )(w-) mixed signs to get - 56
2(w+7)(w-8)=0
Divide both sides by 2
(w+7)(w-8)=0
w+7=0 so w=-7 NO width cannot be negative
w-8=0 so w=8 Yes
Width = 8m
Length = 2(8)-2 = 14m
.......................
Delighted to help.
-Reading Boosters
Wanting for other what we want for ourselves.